5. A consumer consumes Jeans and t-shirts. At his current rate of consumption, his marginal utility of jeans is 60 and his marginal utility of t-shirts is 30. The price of a pair of jeans is $30 and the price of a t-shirt is $10.00. Remember that to maximize utility the consumer should consume such that MUa/Pa=MUb/Pb=MUn/Pn. a. Show the ratio of the marginal utility to price of the two goods. b. Would you suggest he buy more jeans and fewer t-shirts, or more t-shirts and fewer jeans? Explain

6. a Explain the law of diminishing marginal utility.

b. Draw a graph of a total utility curve and connect it with a marginal utility curve. What happens to total utility when marginal utility is zero?

8. a. What are the three types of firms?

b. Explain the advantage and disadvantage of each type of firm.

c. Which type of firm produces the most revenue?

9. a. What are implicit costs?

b. Do implicit cost contribute to the opportunity cost of production? Explain

c. Should an implicit cost be counted as cost in determining profit? Explain

d. Give two examples of implicit costs

e. Does a firm’s accounting statements usually take implicit cost into consideration? Explain

10. Indicate which of the following are long run or short run activities? Don’t just say long run or short run explain your answer.

a. changing the amount of heavy equipment in an auto factory.

b. increasing the number of unskilled workers for a fast food establishment

c. buying more raw materials for a shoe factory

d. increasing the square footage that the cup cake factory occupies

e. training more doctors to improve the health care needs of the inner cities
Answer #1

5. Marginal utility of jeans (MUJ) = 60
Marginal utility of t-shirts (MUT) = 30
Price of jeans, PJ = $30
Price of t-shirts (PT) = $10

a) MUJ/PJ = 60/30 = 2 and MUT/PT = 30/10 = 3
So, utility is not maximized as MUJ/PJ < MUT/PT

b) To maximize utility, the consumer should buy more t-shirts and fewer jeans. This is because by consuming more t-shirts and fewer jeans, MUJ will increase and MUT will decline which will increase MUJ/PJ and decrease MUT/PT until these two ratios become equal and utility is maximized.

(Note: Post one question at a time.)